OpenStudy (anonymous):

Solve for x: the quantity of x plus 20 over 2 = 3x

OpenStudy (bahrom7893):

(x+20)/2=3x x+20=3x*2 x+20=6x x-6x=-20 -5x=-20 x=4

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

First you want to get rid of the denominator so you multiply 2 on both sides. Then you will subtract both sides by x to move it to one side. After which you can continue to isolate x to one side.